OpenStudy (anonymous):

f(x) = -16x2 + 22x + 3 Part A: What are the x-intercepts of the graph of the f(x)? Show your work. (2 points) Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work. (3 points) Part C: What are the steps you would use to graph f(x)? Justify that you can use the answers obtained in Part A and Part B to draw the graph. (5 points)

OpenStudy (igreen):

We know that when y = 0, you get your x intercept, so replace f(x) with 0. \(-16x^2 + 22x + 3 = 0\) Factor: \((-2x + 3)(8x + 1) = 0\) Now we solve each individual one for \(x\) and it will give us our solutions. \(-2x + 3 = 0\) \(-2x = -3\) \(\color{lime}{x = 1.5}\) ----------------------------- \(8x + 1 = 0\) \(8x = -1\) \(\color{lime}{x = -0.125}\) So our \(x\)-intercepts are 1.5 and -0.125.

OpenStudy (igreen):

To find out if the vertex is going to be a minimum or a maximum, plug it into the formula \(c - \dfrac{b^2}{4a}\)

OpenStudy (igreen):

Note: \(ax^2 + bx + c\) \(-16x^2 + 22x + 3\) a = -16 b = 22 c = 3

OpenStudy (igreen):

Plug them into our formula: \(3 - \dfrac{22^2}{4(-16)}\) Simplify: \(3 - \dfrac{484}{-64}\) \(3 + 7.5625\) \(10.5625\) Since the answer is positive, you will have a minimum value.

OpenStudy (igreen):

To find the vertex we use the formula: \(\dfrac{-b}{2a}\). Like I said before: a = -16 b = 22 Plug them in: \(\dfrac{-22}{2(-16)}\) Simplify: \(\dfrac{-22}{-32}\) \(0.6875\) So the x-intercept to the vertex is 0.6875, we can plug this back into our function to get the y-intercept. \(-16x^2 + 22x + 3\) \(-16(0.6875)^2 + 22(0.6875) + 3\) \(-16(0.47265625) + 15.125 + 3\) \(-7.5625 + 15.125 + 3\) \(7.5625 + 3\) \(10.5625\) So the vertex is (0.6875, 10.5625).

OpenStudy (igreen):

You can graph it buy drawing a line through the x-intercepts and the vertex.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @nana1007

OpenStudy (anonymous):

thank you soo much!

OpenStudy (igreen):

No problem!

OpenStudy (igreen):

Btw, the answers are 100% correct. I checked my answers by graphing. https://www.desmos.com/calculator

OpenStudy (igreen):

Yes, you can round them if you want.

OpenStudy (igreen):

Vertex can be rounded to: (0.69, 10.56)